Please show me how to solve for the slope of the line 4x +20Y = 5
Answer by checkley71(8403)

YOU NEED TO PUT THIS EQUATION INTO THE LINE FORMAT
[Y=mX+b] WHERE (m)=SLOPE & (b)=Y INTERCEPT.
4X+20Y=5
20Y=-4X+5
Y=-4X/20+5/20
Y=-X/5+4 THUS THIS LINE HAS A SLOPE(m)=-1/5.